In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate receive?
Answer by faceoff57(108)

x + (x+220) = 810
x + x + 220 = 810
2x + 220 = 810
2x = 810 - 220
2x = 590
x = 295
one candidate received 295 votes.
the second candidate received 515 votes.